the quinhydrone electrode can be used for the potentiometric determination of pH. the solution to be measured is saturated with quinhydrone, an equimolar mixture of quinone and hydroquinone and the potential of the solution is measured with a platinum electrode. the half reaction has an E of 0.699. What is the pH of a solution saturated with quinhydrone if the potential of a platinum electrode in the solution measured against a saturated calomel electrode is -.205V
Answer #1

The electrode reaction is

C6H4O2 (Quinone) + 2H+ + 2e- → C6H4(OH)2 (hydroquinone)

Quinone is oxidant, and hydroquinone is reductant.

For the potentiometric measurements, it is a combination of the quinhydrone electrode with suitable reference electrode to create electrochemical cell. Thus,

pH = (E0(Q/H2Q) – Ecalomel – E) / 2.303 *R*T/F

E0(Q/H2Q) = 0.699 V

E calomel is a standard table data that was supposed to be given to you in a problem conditions along with the Temperature. I suppose it is room T conditions or that would be stated otherwise. Thus, E calomel at T room = 0.2410 V

E is measured = - 0.205 V

T = Room Temperature = 25 oC

R = Gas Constant = 8.314 J kmol-

F = Faraday constant = 96485 C mol-

Plugging in numbers: pH = (0.699 V - 0.2410 V + 0.205 V) / 0.05916 V = 0.663 / 0.05916 = 11.2

Therefore, the pH of a solution saturated = 11.2